If you have a bad grass fire with wind that is burning at the rate of .04 acres per second how many acres will burn in 48 hours? (assuming the rate will not change the entire time)

Answer by ptaylor(2198)

0.04 acres/sec=0.04 *60acres/min= 2.4 acres/min------(60 sec=1 min)
2.4 acres/min=2.4*60 acres/hour=144 acres/hour (60 min=1 hour)
144 acres/hour*48 hours=6912 acres
CK
how many seconds in 1 hour? 60*60=3600
How many seconds in 48 hours? 3600*48=172800
so, 0.04*172800=6912
6912=6912
Hope this helps---ptaylor